OpenStudy (anonymous):

Write a recursive formula for the explicit formula: A(n)=5+(n-1)(3)

OpenStudy (amistre64):

might help to write out the first 3 or 4 terms to see how its moving ... notice a pattern if you will

OpenStudy (anonymous):

What do you mean?

OpenStudy (michele_laino):

hint: we can write: \[\begin{gathered} {a_1} = 5 \hfill \\ {a_2} = 5 + 3 \times (2 - 1) = 5 + 3 = 8 \hfill \\ {a_3} = 5 + 3(3 - 1) = 5 + 6 = 11 \hfill \\ {a_4} = 5 + 3(4 - 1) = 5 + 9 = 14 \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

as you can see, a term is three more of the preceding term

OpenStudy (anonymous):

Okay, yeah I see it

OpenStudy (michele_laino):

so we can write: \[A\left( n \right) = A\left( {n - 1} \right) + ...?\]

OpenStudy (michele_laino):

please complete my formula

OpenStudy (anonymous):

3?

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

Thanks!

OpenStudy (michele_laino):

Thanks!